OpenStudy (anonymous):

Integrate tan^3(x/4)sec^2(x/4)

terenzreignz (terenzreignz):

\[\Large \int \tan^3\left(\frac{x}4\right)\sec^2\left(\frac{x}4\right)dx\] It's like it's just begging for a substitution ^_^

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\tan ^{3}\frac{ x }{ 4 }\sec ^{2}\frac{ x }{ 4 }\]

terenzreignz (terenzreignz):

Mine's better looking... anyway, the presence of a \(\large \sec^2\) which just happens to be the derivative of \(\large \tan\) may indicate that you use a substitution

terenzreignz (terenzreignz):

Letting \(\large u = \tan\left(\frac{x}4\right)\) should do the trick :D

terenzreignz (terenzreignz):

I'm guessing you already have an answer? :D

OpenStudy (anonymous):

Yes, thank you

terenzreignz (terenzreignz):

Good :)